OpenStudy (anonymous):

The area of a circle is 16π cm2. What is the circle's circumference?

OpenStudy (saifoo.khan):

Step 1, find it's radius. Step 2, use that to find the circumference.

OpenStudy (saifoo.khan):

Area = pi*r^2 16 pi = pi r^2 16 = r^2 r = 4 Now, C = 2 pi r C = 2 pi (4)

OpenStudy (anonymous):

The area of a circle is the formula is: A = πr^2 C = 2πr Just work backwards from the area formula and plug it into the circumference formula.